# Distributed Continuous-Time Control via System Level Synthesis

### v3

@@ IV-A Finite-dimensional truncation
Problem (21) involves solving over the infinite-dimensional affine subspace (19). In previous work on discrete-time SLS, finite impulse response approximations are used to truncate this infinite-dimensional problem into a finite-dimensional one. Essentially, instead of solving over all closed-loop responses that satisfy the discrete-time equivalent of (19), we restrict ourselves to closed-loop responses that satisfy (19) and place all closed-loop poles at zero, up to some fixed multiplicity. This technique is not applicable to continuous-time systems. An analogous approach is to first select some fixed, finite multiset $\mathcal{P}$ of closed-loop poles, and restrict ourselves to solving over all closed-loop responses that satisfy (19) and place closed-loop poles at $\mathcal{P}$. This is related to the notion of Ritz approximations, which utilize a single pole with varying multiplicities (i.e., $\mathcal{P}$ contains a single value repeated many times) [12]. It is also possible to use multiple poles [13].
First, we provide a simple result on the truncation of affine subspace (19). Given some multiset of poles $\mathcal{P}$ where all poles are contained in the open left-half-plane (OLHP), we restrict ourselves to considering closed-loop responses of the form:
@@
where $K$ is the number of distinct poles, and $m_{l}$ is the number of times $p_{l}$ appears in $\mathcal{P}$ (i.e., the highest allowable multiplicity for pole $p_{l}$). We will refer to $\Phi_{x}(l,j)$ and $\Phi_{u}(l,j)$ as the spectral components of $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$, respectively. Now, we can rewrite constraint (19) in terms of the spectral components of the closed-loop responses:
@@
By construction, $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$ are strictly causal and stable. It remains to show that (19a) holds. To do so, plug (22) into this constraint and write out the constraint at each index (similar to [14]). ∎
Remark: The number of poles required to satisfy (23) (and therefore (19)) does not necessarily depend on the size of the system. For instance, consider system (1) with an arbitrary number of states and full rank $B$ matrix. Here, it is sufficient to choose a single pole $p<0$. Then, $\mathbf{\Phi}_{x}(s)=\frac{1}{s-p}I_{n}$ and $\mathbf{\Phi}_{u}(s)=\frac{1}{s-p}B^{-1}$ satisfy (23).
@@
We will use simple pole approximation (SPA); all poles have multiplicity 1 (i.e., $m_{l}=1$ $\forall$ $l$). Then, (22) becomes:
$\mathbf{\Phi}_{x}(s)=\sum_{l=1}^{K}\Phi_{x}(l)\frac{1}{(s-p_{l})},\quad\mathbf{\Phi}_{u}(s)=\sum_{l=1}^{K}\Phi_{u}(l)\frac{1}{(s-p_{l})}$ (24)
@@ IV-B Simple pole approximation
We now present the SPA method for continuous-time SLS. Techniques in this section are inspired by those from [13]; our contributions are to devise and explicit continuous-time version, and rigorously demonstrate (Theorems 3-4, Lemma 2) that poles generated via SPA indeed let us approximate the optimal continuous-time closed-loop responses.
Let $\mathcal{P}$ now denote a set of poles chosen via SPA which contains $K$ poles, where $K$ is an even number. Individual poles are denoted by $p_{k}$ where $k\in\{-\frac{K}{2},\dots,-1,1,\dots,\frac{K}{2}\}$. $\mathcal{P}$ is generated as follows:
$\theta_{k}$ $=2\sqrt{\pi k},\quad r_{k}=\sqrt{\frac{k}{K+1}}$ (26)
@@
where we first generate a set of discrete-time poles $z_{k}\in\mathbb{D}$ using the Archimedes spiral and its reflection, then map these poles into the continuous-time OLHP using the bilinear transform. We now show that this approximation method approaches the optimal solution. First, we recall a discrete-time result from prior work.
@@
$\mathcal{Z}_{\mathcal{Q}}$ can represent the poles of any transfer function matrix; thus, this theorem shows that SPA-generated poles can approximate poles of any transfer function matrix in discrete time. We are interested in developing an analogous result for the continuous-time setting. To do so, we first show that if two poles are close in the unit disk, then they must also be close in the OLHP after applying the bilinear transform.
@@
Notice that a right-hand-side term can be upper-bounded:
@@
Plugging this into (28) yields $|p-q|\leq c_{2}|z_{p}-z_{q}|$, the first part of the theorem statement. Combining this with $|z_{p}-z_{q}|\leq c_{1}$ yields $|p-q|\leq c_{1}c_{2}=c_{4}$, the second part of the theorem statement. ∎
This theorem states that if two poles are sufficiently close in $\mathbb{D}$, their bilinear transforms are close in the OLHP. This allows us to analyze SPA for continuous-time poles:
@@
Now, we present the final result: that transfer function matrices using SPA-generated poles can approximate arbitrary stable transfer function matrices in continuous time up to some error bound. We will write transfer function matrices with SPA-generated poles as:
@@
where $\{G_{p}\}_{p\in\mathcal{P}}$ are constant matrices.
@@
Let $\mathcal{P}$ be a set of $K$ poles generated by continuous-time SPA as described in equation (26), where $K$ is selected so that $D(\mathcal{P})\leq c_{4}$. Then, for any transfer function matrix $\mathbf{S}(s)\in\frac{1}{s}\mathcal{RH}_{\infty}$, there exists constants $c_{s},c_{s}^{\prime}>0$ and matrices $\{G_{p}\}_{p\in\mathcal{P}}$ such that
@@
This result is analogous to Theorem 1 in [13]. A proof sketch is provided in the Appendix, where we outline the main ideas and emphasize the key adaptations required for the continuous-time setting.
@@ V $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ Control via SLS
We now describe the full continuous-time SLS formulations for structured $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ control. Instead of working with the infinite-dimensional formulation (21), we will constrain closed-loop responses $\mathbf{\Phi}_{x}$ and $\mathbf{\Phi}_{u}$ to have pre-generated stable poles, and optimize over spectral components. Results can be used with poles generated through SPA or other methods (e.g., Ritz approximation).
To simplify notation, we first order the SPA-generated pole set $\mathcal{P}$ so that complex conjugate pairs are next to one another: $\mathcal{P}=\{p_{1},p_{-1},\dots p_{\frac{K}{2}},p_{-\frac{K}{2}}\}$. For the remainder of the paper (except the Appendix), we re-label these poles as $\mathcal{P}=\{p_{1},p_{2},\dots p_{K}\}$, with the convention that $(p_{2i-1},p_{2i})$ is a complex conjugate pair. While poles can be complex, transfer function matrices $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$ must be real-valued; thus, spectral components must satisfy the following:
$\Phi_{x}(2i-1)$ $=\overline{\Phi_{x}(2i)},$ (34)
$\Phi_{u}(2i-1)$ $=\overline{\Phi_{u}(2i)},\forall i\in\{1,2,\ldots,\frac{K}{2}\}$
Then, for poles chosen via SPA, enforcing feasibility constraints is equivalent to enforcing (34) and (25) on spectral components. Enforcing sparsity constraints (21a) on transfer function matrices is equivalent to enforcing sparsity constraints on their spectral components, i.e.,
@@
Now, we require convex expressions for objective $g$ in terms of spectral components.
@@ V-A State-Space Realization
The objective is to minimize $\|\mathbf{\Psi}(s)\|_{\mathcal{H}_{2}}$ for $\mathcal{H}_{2}$ control, and $\|\mathbf{\Psi}(s)\|_{\mathcal{H}_{\infty}}$ for $\mathcal{H}_{\infty}$ control. We now present a state-space realization of (38) to facilitate this:
@@
When using the SPA-generated poles, all poles are complex, which means realization (V-A) is complex. However, there always exists an orthogonal matrix $T\in\mathbb{R}^{nK\times nK}$ such that the similarity-transformed matrices
@@
are real-valued as long as (34) are satisfied. This can be verified by plugging in
This technique can be generalized to non-repeated, non-SPA-generated poles, so long as $\mathbf{\Psi}(s)$ itself is real. In this case, partition $\mathcal{A}$ into $\mathcal{A}=\text{blkdiag}(\mathcal{A}_{c},\mathcal{A}_{r})$, where $\mathcal{A}_{c}$ contains complex conjugate poles and $\mathcal{A}_{r}$ contains real-valued poles. Poles in $\mathcal{A}_{c}$ should follow the same ordering convention as in the SPA case, i.e., $(p_{2i-1},p_{2i})$ should be a complex conjugate pole pair. Then, apply $T=\text{blkdiag}(U,U,\ldots U,I)$, where the identity matrix is the same size as $\mathcal{A}_{r}$.
In the case of repeated poles (e.g., Ritz), let $\mathbf{\Psi}_{ij}(s)$ denote the $(i,j)^{\text{th}}$ element of transfer function matrix $\mathbf{\Psi}(s)$, with spectral components $\Psi_{ij}(l)$, i.e.,
@@
Here, $\mathcal{A}^{ij}$ is a $K$ by $K$ matrix, and $\mathcal{B}^{ij}$ and $\mathcal{C}^{ij}$ are vectors of length $K$. To find a state-space realization for the full transfer function matrix $\mathbf{\Psi}(s)$, concatenate state-space realizations $(\mathcal{A}^{ij},\mathcal{B}^{ij},\mathcal{C}^{ij})$ across all elements $\mathbf{\Psi}_{ij}(s)$.
The remainder of this section computes $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ objectives using these state-space realizations; they are applicable to SPA- and non-SPA-generated poles. We use ($\tilde{\mathcal{A}},\tilde{\mathcal{B}},\tilde{\mathcal{C}}$) to indicate the appropriate real-valued state-space realization.
@@ V-B $\mathcal{H}_{2}$ Control
The $\mathcal{H}_{2}$ control problem can be formulated via linear matrix inequalities (LMIs) using results from [15]. In particular, $\left\|\mathbf{\Psi}(s)\right\|_{\mathcal{H}_{2}}<\eta$ if $\exists P\succ 0$, $W\succ 0$ such that
@@
Then, we can rewrite problem (36) as:
@@
which is convex in its decision variables. The LMI constraint (44) can be partitioned into smaller parallel LMI constraints by using partitioned matrices (similar idea as (43)); then, this problem is separable and therefore scalable (see Section III). When poles are distinct (e.g., chosen via SPA), the $\mathcal{H}_{2}$ problem can also be written as:
@@ V-C $\mathcal{H}_{\infty}$ Control
The $\mathcal{H}_{\infty}$ control problem can also be formulated via LMIs. We first recall the KYP lemma (Lemma 4.1 in [16]):
@@
Then, we can rewrite problem (36) as:
@@
This problem is convex in its decision variables.
Remark: The size of the $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ problems depend on the number of states and poles. We can reduce the number of poles without sacrificing performance (see Section VI) or sacrificing stability (see remark after Lemma 1).
@@ VI Simulations
We simulate our proposed controller synthesis technique on two plants: a chain and a grid. For all simulations, we set $Q$ and $R$ to be identity matrices of appropriate dimension. Simulation results may be reproduced using the relevant scripts in the SLS-MATLAB toolbox [17].
VI-A Chain Model
We have 11 nodes arranged in a chain; each node is dynamically coupled to its neighbors. $A$ is tri-diagonal, with $A_{ii}=0.6$, $A_{i,i-1}=A_{i-1,i}=0.4$, and $B=I$. This system is unstable. We use an $\mathcal{H}_{2}$ controller with 4 poles and $d=2$; sub-controllers can only communicate with immediate neighbors and their neighbors (see Section II-A). The $\mathcal{H}_{2}$ cost of our SPA-based controller is $1.18$ times the cost of the centralized LQR controller. By design, this controller achieves local communication and local disturbance rejection; the latter can be verified in Figure 3.
Figure 3: Simulation with $d=2$. A disturbance is injected at node 6 at time $t=0$; it propagates only to nodes 4, 5, 6, 7, and 8 (i.e., nodes within $d=2$ distance of node 6) before being rejected.
VI-B Grid Model
We consider linearized swing equations, as is done in [4]. The dynamics of node $i$ on the grid is given by:
$m_{i}\ddot{\theta}_{i}+\beta_{i}\dot{\theta}_{i}=-\sum_{j\in\mathcal{N}_{i}}k_{ij}(\theta_{i}-\theta_{j})+w_{i}+u_{i}$ (50)
where parameters $m_{i}^{-1}$, $\beta_{i}$, and $k_{ij}$ represent inertia, damping, and coupling, respectively; $\theta_{i}$, $w_{i}$, and $u_{i}$ are phase angle deviation, disturbance, and control action, respectively; $\mathcal{N}_{i}$ indicates the immediate neighbors of node $i$. The state of each node can be written as $x]_{i}:={[\theta_{i}\quad\dot{\theta}_{i}]}^{\top}$. Then,
@@ VI-B Grid Model
We generate a random grid topology (shown in Figure 1) and random parameters $m_{i},\beta_{i}$, and $k_{ij}$ using similar distributions as in [4]. The resulting system is marginally stable.
We design a number of $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ controllers with (1) varying degrees of local communication and local disturbance rejection and (2) varying number of poles. Results are shown in Figure 4. As the number of poles or locality parameter $d$ increases, performance becomes more optimal. We remark that this system has 18 states, and the true optimal closed-loop has 18 poles; however, a near-optimal controller can be designed via SPA using only 6 poles, for both $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$. We also include the results for Ritz approximations for $\mathcal{H}_{2}$; we note that for a small number of poles (i.e., low multiplicity), this method is impractical as it is extremely sensitive to the choice of pole placement. We also note that the $\mathcal{H}_{\infty}$ distributed controller experiences far less performance degradation due to reduced poles or reduced communication (via $d$) compared to the $\mathcal{H}_{2}$ controller.
We also compared the performance of our continuous-time SLS controllers to discrete-time SLS controllers with zero-order hold. We let each controller have the same number of free variables; for the discrete-time case, we set the time horizon equal to 7, and for the continuous-time case, we pick 6 poles using the SPA technique. The resulting $\mathcal{H}_{2}$ costs are 5.22 for the continuous-time case, and 5.48, 5.32, and 6.98 for the discrete-time case corresponding to sampling time $\Delta t=0.05$, $0.1$, and $0.2$ seconds respectively. The continuous-time controller outperforms all discrete-time controllers; as expected, there is benefit in designing directly in the continuous-time domain.
Figure 4: SLS for $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ control. Costs are normalized by the optimal centralized controller; this is obtained using lqr for $\mathcal{H}_{2}$ and hinfsyn for $\mathcal{H}_{\infty}$ in MATLAB. (Left) Cost vs. # of poles for centralized (i.e., dense) controllers. For $\mathcal{H}_{2}$ control, we include comparisons between Ritz and SPA methods; for Ritz method, we vary the multiplicity of the pole instead of the number of poles (there is always only one pole). (Center) Cost vs. # of poles for distributed controllers with $d=2$. (Right) Cost vs. $d$ with fixed number of poles (6 poles for $\mathcal{H}_{2}$ and 4 poles for $\mathcal{H}_{\infty}$). For this system, $d=5$ is equivalent to a centralized controller since the diameter of the graph is 5.

### v4

@@ IV-A Finite-dimensional truncation
Problem (11) involves solving over the infinite-dimensional affine subspace (8). In previous work on discrete-time SLS, finite impulse response approximations are used to truncate this problem into a finite-dimensional one; we restrict ourselves to closed-loop responses that satisfy (8) and place all closed-loop poles at zero, up to some fixed multiplicity. For continuous-time systems, an analogous approach is to first select some fixed, finite multiset $\mathcal{P}$ of closed-loop poles, and restrict ourselves to solving over all closed-loop responses that satisfy (8) and place closed-loop poles at $\mathcal{P}$. This is related to the notion of Ritz approximations in continuous time, which utilize a single pole with varying multiplicities [10], as well as simple pole approximation in discrete-time [4].
Given $\mathcal{P}$, some multiset of poles which are contained in the open left-half-plane (OLHP), we restrict ourselves to considering closed-loop responses of the form:
@@
where $K$ is the number of distinct poles in $\mathcal{P}$, and $m_{l}$ is the number of times $p_{l}$ appears in $\mathcal{P}$. We will refer to $\Phi_{x}(l,j)$ and $\Phi_{u}(l,j)$ as the spectral components of $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$, respectively. Now, we can rewrite constraint (8) in terms of the spectral components of the closed-loop responses:
@@
(8b) holds by construction of $\mathcal{P}$ and (12). To show that (8a) holds, plug (12) into (8a) and write out the constraint at each index (similar to [3]). ∎
Remark: The number of poles required to satisfy (13) does not depend on system size. For example, consider system (1) with an arbitrary number of states and full rank $B$ matrix. It is sufficient to choose a single pole $p<0$; then, $\mathbf{\Phi}_{x}(s)=\frac{1}{s-p}I_{n}$ and $\mathbf{\Phi}_{u}(s)=\frac{1}{s-p}B^{-1}$ satisfy (13).
@@
We use simple pole approximation (SPA), where all poles have multiplicity 1 (i.e., $m_{l}=1$ $\forall$ $l$). Then, (12) becomes:
$\mathbf{\Phi}_{x}(s)=\sum_{l=1}^{K}\Phi_{x}(l)\frac{1}{(s-p_{l})},\hskip 5.0pt\mathbf{\Phi}_{u}(s)=\sum_{l=1}^{K}\Phi_{u}(l)\frac{1}{(s-p_{l})}$ (14)
@@ IV-B Simple pole approximation
We now present SPA for continuous-time SLS. Our approach is inspired by [4]; here, we devise a continuous-time version and rigorously demonstrate (Theorems 3-4, Lemma 2) that poles generated via SPA indeed let us approximate the optimal continuous-time closed-loop responses.
Let $\mathcal{P}$ now denote a set of $K$ poles generated by SPA, where $K$ is even. We first generate discrete-time poles $z_{k}\in\mathbb{D}$ using the Archimedes spiral and its reflection, then map them into the OLHP using the bilinear transform:
$\text{for }k=1\ldots\frac{K}{2},\quad\theta_{k}=2\sqrt{\pi k},\quad r_{k}=\sqrt{\frac{k}{K+1}}$ (16)
@@
To show that this approximation method approaches the optimal solution, we first recall a result from prior work.
@@
This theorem shows that SPA-generated poles can approximate the poles ($\mathcal{Z}_{\mathcal{Q}}$) of any discrete-time transfer function matrix. We begin to extend this to continuous-time: first, we show that poles close in the unit disk will also be close in the OLHP after applying the bilinear transform.
@@
One of the right-hand-side terms can be upper-bounded:
@@
Plugging this into (18) yields $|p-q|\leq c_{2}|z_{p}-z_{q}|$. Combining this with $|z_{p}-z_{q}|\leq c_{1}$ yields $|p-q|\leq c_{1}c_{2}=c_{4}$. ∎
This leads to a continuous-time analogue of Theorem 2:
@@
Now, the final result: transfer function matrices using SPA-generated poles can approximate arbitrary stable transfer function matrices in continuous time. We write transfer function matrices with SPA-generated poles as:
@@
where spectral components $\{G_{p}\}_{p\in\mathcal{P}}$ are constant matrices.
@@
Let $\mathcal{P}$ be a set of poles generated by continuous-time SPA per (16); select $K$ so that $D(\mathcal{P})\leq c_{4}$. Then, for any transfer function matrix $\mathbf{S}(s)\in\frac{1}{s}\mathcal{RH}_{\infty}$, there exists constants $c_{s},c_{s}^{\prime}>0$ and matrices $\{G_{p}\}_{p\in\mathcal{P}}$ such that
@@
This result and its proof are analogous to Theorem 1 in [4]; a proof sketch is provided in the Appendix.
@@ V $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ Control via SLS
SLS problem (11) is infinite-dimensional. To bypass this, we constrain closed-loop responses $\mathbf{\Phi}_{x}$ and $\mathbf{\Phi}_{u}$ to have pre-generated stable poles, and optimize over spectral components. These results can be used with poles generated through SPA or other methods (e.g., Ritz).
To simplify notation, we re-order the SPA-generated pole set $\mathcal{P}$: $\mathcal{P}=\{p_{1},p_{-1},\dots p_{\frac{K}{2}},p_{-\frac{K}{2}}\}$. For the remainder of the paper (except the Appendix), we re-label these poles as $\mathcal{P}=\{p_{1},p_{2},\dots p_{K}\}$, with the convention that $(p_{2i-1},p_{2i})$ is a complex conjugate pair. While poles can be complex, transfer function matrices $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$ must be real-valued; thus, spectral components must satisfy the following:
$\quad\forall i\in\{1,2,\ldots,\frac{K}{2}\},\quad\Phi_{x}(2i-1)$ $=\overline{\Phi_{x}(2i)},$ (24)
$\Phi_{u}(2i-1)$ $=\overline{\Phi_{u}(2i)},$
Then, for poles chosen via SPA, feasibility constraints (8) can be enforced via (24) and (15). Similarly, sparsity constraints (10) can be enforced via:
@@
where we optimize over spectral components. We now derive expressions for objective $g$ in terms of spectral components.
@@ V-A State-space realization
The objective is to minimize $\|\mathbf{\Psi}(s)\|_{\mathcal{H}_{2}}$ for $\mathcal{H}_{2}$ control, and $\|\mathbf{\Psi}(s)\|_{\mathcal{H}_{\infty}}$ for $\mathcal{H}_{\infty}$ control. To facilitate this, we use the following state-space realization of (28):
@@
When using the SPA-generated poles, all poles are complex, which means realization (V-A) is complex. However, if (24) is satisfied, then we can construct orthogonal matrix
such that the similarity-transformed matrices
@@
are real-valued. Even when poles are not generated by SPA, as long as $\mathbf{\Psi}(s)$ is real-valued, we can always find a real-valued state-space realization for the controller. For general non-repeated poles, partition $\mathcal{A}$ into $\mathcal{A}=\text{blkdiag}(\mathcal{A}_{c},\mathcal{A}_{r})$, where $\mathcal{A}_{c}$ contains complex conjugate poles (where adjacent values are conjugate pairs, similar to SPA) and $\mathcal{A}_{r}$ contains real-valued poles. Then, apply $T=\text{blkdiag}(U,U,\ldots U,I)$. For repeated poles (e.g., Ritz), let the $(i,j)^{\text{th}}$ element of transfer function matrix $\mathbf{\Psi}(s)$ be written as
@@
where $\mathcal{A}^{ij}\in\mathbb{R}^{K\times K}$. Concatenate state-space realizations $(\mathcal{A}^{ij},\mathcal{B}^{ij},\mathcal{C}^{ij})$ to produce a realization for $\mathbf{\Psi}(s)$.
The remainder of this section computes $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ objectives using these state-space realizations.We use ($\tilde{\mathcal{A}},\tilde{\mathcal{B}},\tilde{\mathcal{C}}$) to indicate the appropriate real-valued state-space realization.
@@ V-B $\mathcal{H}_{2}$ control
From linear matrix inequality (LMI) results in [13], we know that $\left\|\mathbf{\Psi}(s)\right\|_{\mathcal{H}_{2}}<\eta$ if $\exists P\succ 0$, $W\succ 0$ such that
@@
Then, we can write the $\mathcal{H}_{2}$ problem as:
@@
which is convex. Constraint (34) can be partitioned into smaller parallel constraints by using partitioned matrices (similar idea as (33)); then, this problem is separable and therefore scalable (see Section III). When poles are distinct (e.g., generated via SPA), we can equivalently write:
@@ V-C $\mathcal{H}_{\infty}$ control
First, recall the KYP lemma (Lemma 4.1 in [5]):
@@
Using this, we can write the convex $\mathcal{H}_{\infty}$ program:
@@ VI Simulations
We simulate our controller synthesis technique on two plants, setting $Q$ and $R$ as identity matrices. Simulations can be reproduced using the SLS-MATLAB toolbox [9].
VI-A Chain plant
This unstable plant has 11 nodes; $A$ is tri-diagonal, with $A_{ii}=0.6$, $A_{i,i-1}=A_{i-1,i}=0.4$, and $B=I$. We use a distributed $\mathcal{H}_{2}$ controller with 4 SPA-generated poles and locality parameter $d=2$.The $\mathcal{H}_{2}$ cost of our controller is $1.18$ times the cost of the centralized LQR controller. By design, this controller achieves local communication and local disturbance rejection (see Fig. 3).
Figure 3: Chain plant with SLS controller. A disturbance is injected at node 6 at time $t=0$; it propagates only to nodes 4, 5, 6, 7, and 8 (i.e., nodes within $d=2$ distance of node 6) before being rejected.
VI-B Grid plant
We consider linearized swing equations, as is done in [1]. The state of node $i$ is $x]_{i}:={[\theta_{i}\quad\dot{\theta}_{i}]}^{\top}$, with dynamics
@@ VI-B Grid plant
where parameters $m_{i}^{-1}$, $\beta_{i}$, and $k_{ij}$ represent inertia, damping, and coupling, respectively; $\theta_{i}$, $w_{i}$, and $u_{i}$ are phase angle deviation, disturbance, and control action, respectively; $\mathcal{N}_{i}$ indicates neighbors of node $i$. We generate a random grid topology (Fig. 1) and random parameters using similar distributions as in [1]; the resulting plant is marginally stable.
We design $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ controllers with varying locality parameters ($d$) and number of poles. Results are shown in Fig. 4. As $d$ or the number of poles decreases, performance becomes less optimal; the $\mathcal{H}_{\infty}$ controller performance is less sensitive in this regard than $\mathcal{H}_{2}$. Note that near-optimal $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ controllers can be designed via SPA using only 6 poles, even though the true optimal closed-loop has 18 poles. We also include results using Ritz approximations for $\mathcal{H}_{2}$, varying the multiplicity of the pole instead of the number of poles; we find that this method is impractical as it is quite sensitive to the choice of pole placement.
We also compared our continuous-time controllers to discrete-time SLS controllers with zero-order hold. We let each controller have the same number of free variables, using horizon $T=7$ for discrete-time and 6 SPA-generated poles for continuous-time. The resulting $\mathcal{H}_{2}$ costs are 5.22 for the continuous-time case, and 5.48, 5.32, and 6.98 for the discrete-time case corresponding to sampling time $\Delta t=0.05$, $0.1$, and $0.2$ seconds. The continuous-time controller outperforms all discrete-time controllers; as expected, there is benefit in designing directly in the continuous-time domain.
Figure 4: Grid plant with various SLS controllers. All costs are normalized by the cost of the optimal centralized controller. For this system, $d=5$ is equivalent to a centralized controller since the diameter of the graph is 5.